OpenStudy (anonymous):

Solve 3x – 5y = 15 for y.

Parth (parthkohli):

Try subtracting 3x from both sides first ;)

Parth (parthkohli):

\(\Rightarrow -5y = 15 - 3x \) This is what it looks like right now.

OpenStudy (anonymous):

y= -5/3-3

Parth (parthkohli):

Now just divide both sides by -5 :)

Parth (parthkohli):

\(\Rightarrow\Large y = {15 - 3x \over -5} \) \(\Rightarrow \Large y = (15 \div 5) - (3x \div -5)\)

Parth (parthkohli):

\(\Rightarrow y = -3 - \Large {3 \over -5}x\)

Parth (parthkohli):

So that's it

OpenStudy (anonymous):

Could it be your above answer + 15?

OpenStudy (anonymous):

Thats what I've got.

Parth (parthkohli):

+15?

OpenStudy (anonymous):

-3/5x + 15

Parth (parthkohli):

No

OpenStudy (anonymous):

So, what about y = 3/5x -3

OpenStudy (anonymous):

I'm losttttt.

Parth (parthkohli):

Let me see

Parth (parthkohli):

Yep!

OpenStudy (anonymous):

Okay, all you needed to do was take away the negative 3/5

Parth (parthkohli):

Yeah

Parth (parthkohli):

-(-) = +

OpenStudy (anonymous):

Sweet thanks.

OpenStudy (anonymous):

Exactly.

Parth (parthkohli):

You're welcome!

OpenStudy (anonymous):

Couldnt get that out of my brain if I tried.

Parth (parthkohli):

lol